# A Minimalist Approach to LLM Reasoning: from Rejection Sampling to Reinforce

### v1

@@ 1 Introduction
Among these, GRPO stands out as one of the most widely used algorithms for enhancing LLMs on math reasoning tasks due to its success in training DeepSeek-R1 (DeepSeek-AI et al.,, 2025). However, its algorithmic details remain largely undocumented, and it is unclear whether its adoption stems from inherent advantages or, rather, from continuity with methods used in their previous studies. In contrast, RAFT has established itself as one of the simplest and most interpretable baselines, consistently showing good empirical performance in prior literature despite its minimalistic design.
@@
To better understand these dynamics, we conducted ablation studies isolating individual design choices using both Qwen (Yang et al.,, 2024) and LLaMA (Grattafiori et al.,, 2024) models across several Reinforce variants. Our results highlight the following key findings:
@@
These insights highlight the importance of sample selection over algorithmic design in reward-based LLM post-training. The codes of this project are available at https://github.com/RLHFlow/Minimal-RL.
@@ 2 Method
$\mathcal{L}^{\text{RAFT++}}(\theta)=\frac{1}{|\mathcal{D}|}\sum_{x,a\in\mathcal{D}}\frac{1}{|a|}\sum_{t=1}^{|a|}\Big{[}\min\Big{(}s_{t}(\theta),\mathrm{clip}(s_{t}(\theta),1-\epsilon,1+\epsilon)\cdot\mathcal{I}\big{(}r(x,a)=\mathop{\mathrm{argmax}}_{i}r(x,a_{i})\big{)}\Big{)}\Big{]},$ (6)
@@ 3 Experiment Setup
Model Algorithm Math500 Minerva Math Olympiad Bench Average Qwen2.5-Math-7B-base Base 41.3 11.0 18.6 23.6 RAFT 77.4 34.4 37.8 49.9 RAFT++ 80.5 35.8 41.2 52.5 Iterative DPO 75.7 30.5 38.3 48.2 Reinforce 80.6 36.1 42.1 52.9 GRPO 81.6 36.7 43.3 53.9 PPO 79.6 34.8 41.1 51.8 LLaMA-3.2-3B-instruct Base 27.3 8.3 6.5 14.0 RAFT 46.6 16.8 15.5 26.3 RAFT++ 48.8 16.9 16.8 27.5 Reinforce 42.9 14.8 12.5 23.4 GRPO 51.0 18.9 18.4 29.4 PPO 47.7 17.3 16.1 27.0
Table 1: Performance of different algorithms across three benchmarks including Math500 (Hendrycks et al.,, 2021), Minerva Math (Lewkowycz et al.,, 2022), and Olympiad Bench (He et al.,, 2024). The reported accuracy is average@16 with a temperature 1.0 and a maximal generation length of 4096 tokens.
@@ 4 Main Results
We summarize the test accuracy of models trained using various algorithms in Table 1. Our first observation is that RAFT (and its variant RAFT++), which is arguably the simplest algorithm, achieves competitive performance compared to more complex methods such as iterative DPO and deep RL-based approaches. Specifically, with Qwen2.5-Math-7B-base, vanilla RAFT reaches an average accuracy of 49.9%, outperforming iterative DPO (48.2%) and approaching PPO (51.8%). With the additional importance sampling and clipping techniques, RAFT++ further improves over vanilla RAFT, achieving 52.5% average accuracy. This result is remarkably close to the state-of-the-art deep RL method GRPO, which reaches an average accuracy of 53.9% in its best model. A similar trend is observed on the LLaMA-3.2-3B-instruct model, demonstrating the robustness of RAFT and RAFT++ across different models. These results are somewhat counter-intuitive, as RL methods are often believed to be more powerful due to their ability to utilize negative feedback. Interestingly, in the LLaMA-based setting, Reinforce performs substantially worse than RAFT++, with an average accuracy of 23.4% compared to the 27.5% of RAFT++. One possible explanation is that defining negative samples solely based on final answer correctness may be too coarse, potentially limiting the benefits of using negative feedback in RL training. We will also include more ablation studies to investigate the role of the negative samples in current practice of RL training.
@@ 4.1 Ablation Study
These findings suggest that negative samples play a crucial role in maintaining exploration and preventing distributional collapse. This exploration benefit is likely a contributing factor to the performance gap between RAFT++ and RL-based methods such as Reinforce and GRPO. To further investigate the relationship between policy entropy and reward learning, we incorporate the “clip higher” technique from Yu et al., (2025), which uses an asymmetric clipping range with $\epsilon_{1}=0.1$ for the lower bound and a larger $\epsilon_{2}=0.2$ for the upper bound. We apply this variant to the LLaMA-3.2-3B-instruct model and visualize both the training reward curves and policy entropy curves in the right figure of Figure 2. Consistent with the findings in Yu et al., (2025), using a larger $\epsilon_{2}$ helps stabilize the policy entropy over the online training. As a result, this enhanced RAFT++ variant outperforms the original RAFT++ during the later stages of training.

### v2

@@ 1 Introduction
Among these, GRPO stands out as one of the most widely used algorithms for enhancing LLMs on math reasoning tasks due to its success in training DeepSeek-R1 (DeepSeek-AI et al.,, 2025). However, its algorithmic details remain largely undocumented, and it is unclear whether its adoption stems from inherent advantages or, rather, from continuity with methods used in their previous studies. To the best of our knowledge, a comprehensive justification of the algorithmic advantage of GRPO is still missing so far. In contrast, RAFT has established itself as one of the simplest and most interpretable baselines, consistently showing good empirical performance in prior literature despite its minimalistic design.
@@
To better understand these dynamics, we conducted ablation studies isolating individual design choices using both Qwen (Yang et al., 2024a, ) and LLaMA (Grattafiori et al.,, 2024) models across several Reinforce variants. Our results highlight the following key findings:
@@
These insights highlight the importance of sample selection over algorithmic design in reward-based LLM post-training. The codes of this project will be publicly available with detailed training scripts. The codes of this project are available at https://github.com/RLHFlow/Minimal-RL.
2 Related Works
Most of the related works on RL algorithm design for LLMs have been discussed in the introduction. Here, we review a few that are most relevant to our project.
Data filtering in LLM Post-Training.
Several recent works in RLHF and preference optimization explore data filtering strategies to improve training quality. For example, Yuan et al., (2024); Dong et al., (2024); Xiong et al., (2024); Shen and Zhang, (2024) discard the candidates except for the top and bottom-ranked responses to reduce noise in pairwise comparisons during RLHF learning. Yu et al., 2025a further incorporates reward and length information of rejected responses into the filtering process. For reasoning tasks, it is also common to remove prompts that are too easy or too hard (Yang et al., 2024b, ; Zhao et al.,, 2024), though this is typically done once before training. In contrast, our proposed Reinforce-Rej performs filtering online throughout training. Furthermore, our study reveals a connection between the strong empirical performance of GRPO and implicit data filtering. Reinforce-Rej can be seen as a natural extension of these insights from our ablation studies.
LLM for Mathematical Reasoning.
LLMs designed for (mathematical) reasoning have received significant attention, especially following the release of GPT-o1 by OpenAI (Jaech et al.,, 2024) and DeepSeek-R1 by DeepSeek (DeepSeek-AI et al.,, 2025). Earlier efforts primarily focused on building synthetic datasets and applying supervised fine-tuning (Gou et al.,, 2023; Yue et al.,, 2023; Yu et al.,, 2023; Toshniwal et al.,, 2024). In contrast, these new models (o1 and R1) adopt RL with verifier-based rewards as their main training approach. A key difference is that models like GPT-o1 and DeepSeek-R1 use more complex reasoning strategies—such as backward search and self-correction—and tend to generate longer outputs at inference time for better performance. Their success has inspired a surge of open-source efforts to replicate or adapt these training strategies to other domains (Jin et al.,, 2025; Xiong et al.,, 2025; Wang et al.,, 2024). Notably, GRPO has become the default RL method in many of these projects, often without justification. However, whether GRPO is truly better than Reinforce, and (if the answer to the first question is yes) what contributes to its performance gains, remains largely under-explored.
@@ 3 Method
$\mathcal{L}^{\text{RAFT++}}(\theta)$ $=\frac{1}{|\mathcal{D}|}\sum_{x,a\in\mathcal{D}}\frac{1}{|a|}\sum_{t=1}^{|a|}\Big{[}\min\Big{(}s_{t}(\theta),\mathrm{clip}(s_{t}(\theta),1-\epsilon,1+\epsilon)\Big{)}\mathcal{I}\big{(}r(x,a)=\mathop{\mathrm{argmax}}_{i}r(x,a_{i})\big{)}\Big{]},$ (6)
@@ 4 Experiment Setup
Model Algorithm Math500 Minerva Math Olympiad Bench Average Qwen2.5-Math-7B-base Base 41.3 11.0 18.6 23.6 RAFT 77.4 40.8 38.6 52.3 RAFT++ 80.2 44.9 43.3 56.1 Iterative DPO 76.0 31.2 39.3 48.8 Reinforce 80.1 40.7 40.9 53.9 GRPO 81.3 45.5 42.2 56.3 PPO 79.0 39.3 39.1 52.5 Reinforce-Rej 81.9 44.2 43.1 56.4 LLaMA-3.2-3B-instruct Base 26.3 7.4 5.5 13.1 RAFT 46.1 17.6 13.9 25.9 RAFT++ 47.4 19.1 16.3 27.6 Reinforce 45.9 13.7 13.0 24.2 GRPO 49.2 19.3 16.8 28.4 PPO 46.5 19 15.1 26.9 Reinforce-Rej 50.1 19.3 16.1 28.5
Table 1: Performance of different algorithms across three benchmarks including Math500 (Hendrycks et al.,, 2021), Minerva Math (Lewkowycz et al.,, 2022), and Olympiad Bench (He et al.,, 2024). We tune all the algorithms to their best performance by fully optimizing the hyper-parameters (including batch size, mini batch size, and actor learning rate). See Appendix for the detailed parameter setup. The reported accuracy is average@16 with a temperature 1.0 and a maximal generation length of 4096 tokens.
@@ 5 Main Results
We summarize the test accuracy of models trained using various algorithms in Table 1. Our first observation is that RAFT (and its variant RAFT++), which is arguably the simplest algorithm, achieves competitive performance compared to more complex methods such as iterative DPO and deep RL-based approaches. Specifically, with Qwen2.5-Math-7B-base, vanilla RAFT reaches an average accuracy of 52.3%, outperforming iterative DPO (48.8%) and approaching PPO (52.5%). With the additional importance sampling and clipping techniques, RAFT++ further improves over vanilla RAFT, achieving 56.1% average accuracy. This result is remarkably close to the state-of-the-art deep RL method GRPO, which reaches an average accuracy of 56.3% in its best model. A similar trend is observed on the LLaMA-3.2-3B-instruct model, demonstrating the robustness of RAFT and RAFT++ across different models. These results are somewhat counter-intuitive, as RL methods are often believed to be more powerful due to their ability to utilize negative feedback. Interestingly, in the LLaMA-based setting, Reinforce performs substantially worse than RAFT++, with an average accuracy of 24.2% compared to the 27.6% of RAFT++.
One possible explanation is that defining negative samples solely based on final answer correctness may be too coarse, potentially limiting the benefits of negative signals. Moreover, Reinforce with binary reward (±$\pm 1$) can be viewed as fine-tuning on the positive samples and unlearning on the negative samples. When the negative signals are not sufficiently fine-grained, unlearning on the negative samples is more unstable than fine-tuning on the positive samples. We will also include more ablation studies to investigate the role of the negative samples in current practice of RL training.
@@ 5.1 Ablation Study
These findings suggest that negative samples play a crucial role in maintaining exploration and preventing distributional collapse. This exploration benefit is likely a contributing factor to the performance gap between RAFT++ and RL-based methods such as Reinforce and GRPO. To further investigate the relationship between policy entropy and reward learning, we incorporate the “clip higher” technique from Yu et al., 2025b , which uses an asymmetric clipping range with $\epsilon_{1}=0.2$ for the lower bound and a larger $\epsilon_{2}=0.28$ for the upper bound. We apply this variant to the LLaMA-3.2-3B-instruct model and visualize both the training reward curves and policy entropy curves in the right figure of Figure 2. Consistent with the findings in Yu et al., 2025b , using a larger $\epsilon_{2}$ helps stabilize the policy entropy over the online training. As a result, this enhanced RAFT++ variant outperforms the original RAFT++ during the later stages of training.